OpenStudy (mollykenney):

Solve: -4(2x - 1) = -6(x + 2) - 2

OpenStudy (arabpride):

Do you know how to use the distributive property?

OpenStudy (anonymous):

you would distribute which means you'd multiply -4 by 2x and by -1 then -6 by x and 2.

OpenStudy (mollykenney):

yes

OpenStudy (mollykenney):

so -8x+ 4= -6x - 12 - 2

OpenStudy (anonymous):

yes. then you would combine like terms on each side of the equation

OpenStudy (arabpride):

Like this: -4(2x - 1) = -6(x + 2) - 2 (-4 * 2x) (-4* -1) = (-6 * x) (-6 *+2) - 2

OpenStudy (mollykenney):

ok so -8x +4 = -6x - 14

OpenStudy (mollykenney):

18 = 2x = 9?

OpenStudy (arabpride):

-4(2x - 1) = -6(x + 2) - 2 -8x + 4 = -6x - 12 - 2 And combine like terms: -8x + 4 = -6x - 12 - 2 +8x +8x 4 = 2x -12 - 2 (-12 - 2) 4 = 2x - 14 +14 +14 18 = 2x /2 /2 x = 9

OpenStudy (arabpride):

Yep :D x DOES equal 9

OpenStudy (anonymous):

yes :)